# Plancraft: an evaluation dataset for planning with LLM agents

### v1

@@ Abstract
We present Plancraft, a multi-modal evaluation dataset for LLM agents. Plancraft has both a text-only and multi-modal interface, based on the Minecraft crafting GUI. We include the Minecraft Wiki to evaluate tool use and Retrieval Augmented Generation (RAG), as well as an oracle planner and oracle RAG information extractor, to ablate the different components of a modern agent architecture. To evaluate decision-making, Plancraft also includes a subset of examples that are intentionally unsolvable, providing a realistic challenge that requires the agent not only to complete tasks but also to decide whether they are solvable at all. We benchmark both open-source and closed-source LLMs and strategies on our task and compare their performance to a handcrafted planner. We find that LLMs and VLMs struggle with the planning problems that Plancraft introduces, and we offer suggestions on how to improve their capabilities.
@@ 1 Introduction
With the increased performance and affordability of Large Language Models (LLMs) (OpenAI, 2024; Dubey et al., 2024), LLM-based agents have exploded in popularity (Xi et al., 2023). We define LLM agents as systems that incorporate at least one call to an LLM, which influences its autonomous decisions about actions in an environment. LLMs are promising as backbones for agent-based systems because they can leverage general knowledge acquired during pre-training to tackle tasks expressed in natural language. This enables human interactions during planning and execution, which are challenging for other planning paradigms such as Reinforcement Learning (RL) or symbolic planners (Xi et al., 2023).
LLMs can query and interpret knowledge bases, ask questions, and incorporate new information via text, all through a flexible dialogue interface. Even though they’re not designed as planners, they are increasingly applied to embodied environments where various methods and strategies (Yao et al., 2023b; Shinn et al., 2023; Yao et al., 2023a) aim to maximise the performance of autonomous agents (Xi et al., 2024). But LLMs still exhibit problematic issues that limit their reliability and usefulness for general-purpose agents – hallucinations and brittleness to inputs, limited context windows, and lack of grounding when placed in new environments.
@@
Table 1: Comparison of different interactive datasets commonly used in agentic LLM evaluation. Each of these datasets expresses the goal or task in natural language and requires multi-step planning. We design Plancraft to provide a testing ground for which there exists both a natural language knowledge base and a handcrafted planner. Plancraft also includes a subset of the data that is intentionally unsatisfiable (Impossible Set), since the capacity to predict that there’s no valid plan is important. GAIA allows tool use, including web search, however a knowledge base on how to solve the task is not provided.
Most existing benchmarks for evaluating LLM agents focus almost exclusively on success rates (Shridhar et al., 2021; Yao et al., 2022; Liu et al., 2023; Xi et al., 2024): that is, the proportion of trials in which the agent achieved a goal state. Success rates measure whether the agent constructed a valid plan, but do not measure the efficiency or the quality of that plan. Some environments allow an arbitrary number of incorrect steps, while others immediately terminate an episode if a wrong action is executed. Reporting or measuring only success rates also introduces a dataset bias, as it implies that the solution to a given problem is easily verifiable, yet difficult to obtain. As a result, we argue that success rate alone is insufficient to capture the complexity of real-world scenarios. Each additional step of inference incurs non-negligible costs, so metrics should include more fine-grained assessments, such as how close the LLM’s plan is to a handcrafted solution. Furthermore, effective agent-based systems should recognise when a task is unsolvable, as many real-world tasks may lie beyond the agent’s capabilities and indeed many real-world tasks might be difficult to verify as doable, or not. Without the capacity to predict if there’s no valid plan, an agent will incur significant costs in continually monitoring and replanning.
@@
We evaluate different LLM-based agents on our dataset, thereby providing LLM-agent baselines. We test both open-source and closed-source models and evaluate them against different sets of possible actions. For open-source models, we evaluate the impact of fine-tuning agents on a set of expert plans, and compare multi-modal agents to text-only models. We also fine-tune a bounding-box detection model on our environment to provide an interface through which text-only LLMs can interact with the multi-modal environment. We release all our baseline models and code along with the dataset and environment as a stand-alone Python package. [^1] https://github.com/gautierdag/plancraft
@@ 2.1 LLM Agents
Various strategies have been proposed to leverage LLMs as agents with interactive feedback (Yao et al., 2023b; Huang et al., 2022; Shinn et al., 2023; Yao et al., 2023a; Wang et al., 2022b). The basic idea is to harvest LLMs for information, which, if novel to the agent, may result in better behaviours than they would perform otherwise. The simplest strategy, ReAct (Yao et al., 2023b), consists of interleaving actions with ‘thinking’ steps where the LLM is allowed unconstrained generation. Other broader strategies, such as Reflexion (Shinn et al., 2023) or Inner Monologue (Huang et al., 2022) try to promote self-corrective behaviour through external modules or steps. Self-Consistency (Wang et al., 2022b) and Tree-of-Thought (Yao et al., 2023a) sample the space of possible paths and select a solution through a majority vote or through another LLM evaluation step. Although all of these techniques are simple to implement, they can often add significant overhead in inference compute.
@@
One limitation of LLMs is the size of their context window. To address this, modern systems often incorporate a form of Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), which searches a larger set of documents and then restricts the context to only relevant examples (by some quantitative metric), given the current observations and task. For example, Voyager (Wang et al., 2023a) augments its context with the most likely applicable skills for a given problem. Similarly, JARVIS-1 (Wang et al., 2023b) stores successful trajectories and augments the context with a relevant subset, estimated via similarity with the current observation. Since crafting is a core part of Minecraft, its Wiki contains a page with details of recipes for crafting all items. The Minecraft Wiki is therefore well suited to a RAG pipeline where an agent can query the Wiki as a knowledge source.
@@
Enabling LLMs to use external tools, such as web search, calculators, and database queries, can also significantly expand their capabilities beyond simple next-token prediction. This has inspired numerous research efforts (Nakano et al., 2022; Schick et al., 2023; Parisi et al., 2022; Yang et al., 2023; Patil et al., 2023) to integrate and evaluate tool use (or external actions) into LLMs. Some datasets, such as GAIA (Mialon et al., 2023), even evaluate agents without any restriction on the tools allowed. In the context of LLM agents, actions can be seen as analogous to tools.
2.2 LLM Evaluation Datasets
@@
Research on LLMs-based agents requires datasets and benchmarks to evaluate them. These datasets vary significantly in the types of environments in which actions are executed, ranging from simplified text-only worlds to multi-modal environments. ALFWorld (Shridhar et al., 2021) and ALFRED (Shridhar et al., 2020) provide virtual home environments where the agent must manipulate objects to fulfil the instruction, expressed in natural language. Mind2Web (Deng et al., 2023), WebShop (Yao et al., 2022), and WebArena (Zhou et al., 2023) evaluate agents in web-based environments, with tasks requiring interaction with browser interfaces or e-commerce platforms. BabyAI (Chevalier-Boisvert et al., 2019) is a 2D maze-solving game with natural language instructions. Minecraft is also a popular benchmark for agent evaluation, with various datasets such as MineRL (Guss et al., 2019), MineDojo (Fan et al., 2022), and TextCraft Prasad et al. (2023). In Table 1, we compare Plancraft against popular LLM evaluation datasets. While these vary in domain and implementation, all require agents to do multi-step planning to achieve tasks expressed in natural language.
Knowledge Base
@@
Minecraft has a unique advantage in that it is a real game, rather than a construction designed for agent evaluation, and therefore there exists a variety of online content to assist human players. Several works have taken advantage of knowledge bases in Minecraft in conjunction with the game environment. Baker et al. (2022) collected a large dataset of Minecraft videos and trained a model, Video Pretraining (VPT), to solve tasks directly from pixels in MineRL (Guss et al., 2019). Fan et al. (2022) introduced MineDojo along with a knowledge base of scraped resources such as Minecraft videos and pages taken from the Minecraft Wiki and Reddit forum; however, they do not evaluate whether these improve performance. In Plancraft, we collect the pages of recipes available on the Minecraft Wiki and use them to build a knowledge base to evaluate RAG capabilities. Additionally, we implement an oracle recipe search to provide an upper estimate of a well-performing RAG system.
Planner
@@
Although all of the datasets listed in Table 1 can evaluate whether an agent achieves the goal, these lack a handcrafted policy to evaluate the quality of the agent’s plans: they only measure whether the agent has reached the goal. Of the datasets that we compare to, only BabyAI (Chevalier-Boisvert et al., 2019) and ALFRED (Shridhar et al., 2020) release handcrafted policies or solvers. BabyAI implements a Bot Agent with handcrafted policies that can act as a teacher agent for learners in the environment. ALFRED releases expert demonstrations found using a PDDL solver and reports Path Weighted Metrics that take the length of the expert into account. Similarly, we implement a planner for Plancraft to provide a benchmark against which to compare agents.
Impossible Set
All existing datasets lack a mechanism for assessing task feasibility or for handling scenarios where tasks are fundamentally unsolvable, which introduces a bias toward tasks that always assume a solution exists. Plancraft addresses this gap by including an intentional set of impossible tasks – tasks that cannot be solved under the provided constraints. By incorporating these examples, we assess an agent’s capacity to reason about task feasibility. This setup promotes more realistic and cost-effective decision making, encouraging agents to balance problem solving with the practicality of task solvability.
@@ 3 Plancraft
To create Plancraft, we implement the logic and visual representation of the Minecraft crafting GUI. Similarly to Prasad et al. (2023), we reconstruct the crafting process entirely in Python using game files and images of items obtained from the Wiki. Using image manipulation to overlay items on top of the inventory background significantly improves performance, as the observations do not require the Java game engine. Our environment has a one-to-one pixel mapping with the real Minecraft interface and supports multiple resolutions.
@@ 3.1 Assumptions
Plancraft makes a number of simplifying assumptions:
  1. 1.
    Single Agent: there is a single agent in the environment and any changes to the environment are a direct result of this agent’s actions.
  2. 2.
    Deterministic: valid actions are always executed and their effects are predictable.
  3. 3.
    Observable: the crafting interface and inventory are fully observable, either through the text description or through the multi-modal image input.
  4. 4.
    Sequential: aligned with classical planning, each action is discrete and executed one at a time.
@@ 3.2 Action and Observation Space
The abstraction level chosen for the action and observation space has a great impact on the tractability of the planning problem. Other Minecraft environments such as MineDojo (Fan et al., 2022) and Textcraft (Prasad et al., 2023) provide a high-level ‘craft’ command that doesn’t require the agent to manipulate the underlying ingredients. Plancraft, on the other hand, requires the correct placement of ingredients in the inventory and, as such, provides a lower-level symbolic action space. The two possible environment actions are smelt and move, and both actions require a slot_from, a slot_to and a quantity. This abstracts control dynamics and focuses on planning and spatial reasoning.
@@
In the Minecraft crafting GUI, there are $46$ possible item slots or positions. The first of these is the crafting slot and is only populated with an item if the correct items have been deposited within the $3\times 3$ crafting grid. Items can only be withdrawn from the crafting slot, and moving the item out of the slot confirms the crafting operation and removes the items present in the crafting grid. The crafting grid is an arrangement of slots $3\times 3$, where items can be moved in and out. Crafting recipes in Minecraft are either Shapeless or Shaped. Shapeless recipes only check whether the items required are present in the crafting grid, whereas Shaped recipes require the items to also be placed in the correct positions. When the items on the crafting grid match one of the $634$ recipes, then the resulting item is added to the crafting slot. The remaining $36$ slots are generic inventory slots where the player can store items. We denote the crafting slot as [0], the $3\times 3$ crafting grid as slots [A1] to [C3] where the letter denotes each row and the number denotes the column, and the remaining inventory slots as slots [I1] to [I36]. See Figure 2 for an example of a text description of the inventory that uses our annotation scheme.
Plancraft provides two types of observations:
  • text-only: Agents receive a text description of the current inventory and crafting goal. The inventory is perfectly represented using our annotation scheme.
  • multi-modal: Agents receive an image (high resolution with dimensions $664\times 704$) that perfectly contains the crafting GUI. The goal is still provided as text.
@@ 3.3 Task Setup
To generate crafting tasks, we represent the dependencies between items as a tree where each item is crafted from a set of required resources. The target item is the root of the tree, while the leaves and intermediate nodes represent the materials required to craft the item. For each crafting task, we sample the required recipes for the target item and then recursively explore the necessary materials to craft it. This builds a step-by-step plan from raw materials to the final product, simulating the crafting processes with varying levels of complexity. In addition to the necessary crafting materials, we also sample a number ($4$, $8$, $16$) of distractor items to make identifying the correct crafting path more challenging.
Finally, we validate that all generated tasks can be solved in the text-only and multi-modal environments using a planner (see Section 3.4) and exclude any invalid examples. We exclude any paths where the handcrafted planner’s trajectory exceeds 30 steps.

### v2

@@ Abstract
We present Plancraft, a multi-modal evaluation dataset for LLM agents. Plancraft has both a text-only and a multi-modal interface, based on the Minecraft crafting GUI. We include the Minecraft Wiki to evaluate tool use and Retrieval Augmented Generation (RAG), as well as a handcrafted planner and Oracle Retriever, to ablate the different components of a modern agent architecture. To evaluate decision-making, Plancraft also includes a subset of examples that are intentionally unsolvable, providing a realistic challenge that requires the agent not only to complete tasks but also to decide whether they are solvable at all. We benchmark both open-source and closed-source LLMs and compare their performance and efficiency to a handcrafted planner. Overall, we find that LLMs and VLMs struggle with the planning problems that Plancraft introduces, and offer suggestions on how to improve their capabilities.
@@ 1 Introduction
With the increased performance and affordability of Large Language Models (LLMs) (OpenAI, 2024; Dubey et al., 2024), LLM-based agents have exploded in popularity (Xi et al., 2023). We define LLM agents as systems that incorporate at least one call to an LLM which influences its autonomous decisions about actions in an environment. LLMs are promising for agent-based systems because they can leverage general knowledge acquired during pretraining to tackle tasks expressed in natural language. They can also query and interpret knowledge bases, ask questions, and incorporate new information via text or images, all through a flexible dialogue interface. This potential for human interaction during planning and execution is challenging for other planning paradigms such as Reinforcement Learning (RL) or symbolic planners (Xi et al., 2023). Various methods and strategies (Yao et al., 2023a; b; Shinn et al., 2023) aim to maximise the performance of LLM-based autonomous agents (Xi et al., 2024). But LLMs still exhibit problematic issues that limit their reliability and usefulness for planning—hallucinations and brittleness to inputs, limited context windows, and lack of grounding when placed in new environments.
@@
Table 1: A comparison of different interactive datasets commonly used in evaluating LLM agents. Each dataset expresses goals or tasks in natural language and requires multi-step planning. We design Plancraft to provide a testing ground for which there exists both a natural language knowledge base and a handcrafted planner. Plancraft also includes data that is intentionally unsatisfiable (the Impossible Set), since the capacity to predict that there’s no valid plan is important. GAIA allows tool use, including web search, however a knowledge base on how to solve the task is not provided.
Most existing benchmarks for evaluating LLM agents in interactive environments focus on success rates (Shridhar et al., 2021; Yao et al., 2022; Liu et al., 2023; Xi et al., 2024): that is, the proportion of trials in which the agent achieves a goal state. Success rates measure whether the agent constructed a valid plan, but do not measure efficiency or the plan’s quality. Some environments allow an arbitrary number of incorrect steps, while others immediately terminate an episode if a wrong action is executed. Reporting or measuring only success rates also introduces a dataset bias, as it implies that the solution to a given problem is easily verifiable, yet difficult to obtain. As a result, we argue that success rate alone is insufficient to capture the complexity of real-world scenarios. Each additional step of inference incurs non-negligible costs, so metrics should include more fine-grained assessments, such as how close the LLM’s plan is to a handcrafted solution. Furthermore, effective agent-based systems should recognise when a task is unsolvable, as many real-world tasks may lie beyond the agent’s capabilities and indeed many real-world tasks might be difficult to verify as doable, or not. Without the capacity to predict if there’s no valid plan, an agent will incur significant costs in continually monitoring and replanning.
@@
We evaluate different LLM-based agents on our dataset, thereby providing LLM-agent baselines. We test both open-source and closed-source models and evaluate them against different sets of possible actions. For open-source models, we evaluate the impact of fine-tuning agents on a set of expert plans, and compare multi-modal vs. text-only agents. We also fine-tune a bounding-box detection model on our environment to provide an interface through which text-only LLMs can interact with the multi-modal environment. Finally, we release all baseline models and code along with the dataset and environment as a stand-alone Python package. [^1] https://github.com/gautierdag/plancraft
@@ 2.1 LLM Agents
Various strategies have been proposed to leverage LLMs as agents with interactive feedback (Huang et al., 2022; Shinn et al., 2023; Yao et al., 2023a; b; Wang et al., 2022b). The basic idea is to harvest LLMs for information, which, if novel to the agent, may result in better behaviours than they would perform otherwise. The simplest strategy, ReAct (Yao et al., 2023b), consists of interleaving actions with ‘thinking’ steps where the LLM is allowed unconstrained generation. Other broader strategies, such as Reflexion (Shinn et al., 2023) or Inner Monologue (Huang et al., 2022) try to promote self-corrective behaviour through external modules or steps. Self-Consistency (Wang et al., 2022b) and Tree-of-Thought (Yao et al., 2023a) sample the space of possible paths and select a solution through a majority vote or through another LLM evaluation step. Although all of these techniques are simple to implement, they often add significant overhead in inference compute.
@@
One limitation of LLMs is the size of their context window. This is often addressed with Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), which restricts the context to only relevant examples (by some quantitative metric) from the larger dataset, given the current observations and task. For example, contexts in Voyager (Wang et al., 2023a) consist of the most likely applicable skills for a given problem; those in JARVIS-1 (Wang et al., 2023b) are successful trajectories that are sufficiently similar to current observations. Since crafting is a core part of Minecraft, its Wiki contains a page with details of recipes for crafting all items. The Minecraft Wiki is therefore well suited to a RAG pipeline where an agent can query the Wiki as a knowledge source.
@@
Enabling LLMs to use external tools, such as web search, calculators, and database queries, can also significantly expand their capabilities. This has inspired numerous research efforts (Nakano et al., 2022; Schick et al., 2023; Parisi et al., 2022; Yang et al., 2023; Patil et al., 2023) to integrate and evaluate tool use (or external actions) into LLM agents. Some datasets, such as GAIA (Mialon et al., 2023), even evaluate agents without any restriction on the tools allowed.
Beyond RAG and tool use, some work has examined the limitations of LLMs in planning tasks. Kambhampati et al. (2024) argue that while LLMs struggle with long-term planning, they can still complement traditional methods in an LLM-modulo framework. Several datasets like PlanBench (Valmeekam et al., 2023), TRAC (He et al., 2023), and ACPBench (Kokel et al., 2025) were proposed to evaluate planning-related reasoning. ACPBench and TRAC primarily test planning capabilities through multiple-choice or boolean questions, while PlanBench uses LLMs to generate full plans directly given different testing conditions. However, since these datasets do not involve interactive execution, we do not include them in Table 1.
2.2 Interactive LLM Evaluation Datasets
@@
Research on LLM agents requires datasets and benchmarks for evaluation. These datasets vary significantly in the types of environments in which actions are executed, ranging from simplified text-only worlds to multi-modal environments. ALFWorld (Shridhar et al., 2021) and ALFRED (Shridhar et al., 2020) provide virtual home environments in which the agent must manipulate objects to fulfil the instruction, expressed in natural language. Mind2Web (Deng et al., 2023), WebShop (Yao et al., 2022), and WebArena (Zhou et al., 2023) evaluate agents in web-based environments, with tasks requiring interaction with browser interfaces or e-commerce platforms. BabyAI (Chevalier-Boisvert et al., 2019) is a 2D maze-solving game with natural language instructions. Minecraft is also a popular benchmark for agent evaluation, with various datasets such as MineRL (Guss et al., 2019), MineDojo (Fan et al., 2022), and TextCraft (Prasad et al., 2023). In Table 1, we compare Plancraft against these LLM evaluation datasets. While these vary in domain and implementation, they all require agents to do multi-step planning in interactive environments to achieve tasks expressed in natural language.
@@
Knowledge Base.  Minecraft has the advantage that it is a game humans play, rather than constructed only for agent evaluation. So LLM agents can take advantage of existing online content that assists human players. Baker et al. (2022) collected a large dataset of Minecraft videos and trained a model, Video Pretraining (VPT), to solve tasks directly from pixels in MineRL (Guss et al., 2019). Fan et al. (2022) introduced MineDojo along with a knowledge base of scraped resources such as Minecraft videos and pages taken from the Minecraft Wiki and Reddit forum; however, they do not evaluate whether these improve performance. In Plancraft, we collect the pages of recipes on the Minecraft Wiki and use them to build a knowledge base to evaluate RAG capabilities. We also implement an Oracle Retriever for recipes to provide an upper estimate of a well-performing RAG system.
@@
Planner.  All of the datasets listed in Table 1 can evaluate whether an agent achieves the goal. But evaluating a plan’s quality requires a designated optimal policy for comparison. Of the datasets that we compare to, only BabyAI (Chevalier-Boisvert et al., 2019) and ALFRED (Shridhar et al., 2020) release handcrafted policies or solvers. BabyAI implements a Bot Agent with handcrafted policies that can act as a teacher agent for learners in the environment. ALFRED releases expert demonstrations found using a PDDL solver and reports Path Weighted Metrics that take the length of the expert into account. Similarly, we implement a planner for Plancraft to provide a benchmark against which to compare agents (Section 3.4). While PDDL-based classical planning approaches (McDermott et al., 1998) could theoretically be applied to Plancraft, the domain presents unique challenges due to large search space (see Appendix D.1).
Impossible Set.  Most interactive datasets assume that every task has a solution. So they cannot evaluate mechanisms that assess feasibility or handle inherently unsolvable tasks. This can be particularly costly for strategies like Tree-of-Thought, which expand the search space by extensively sampling the LLM for potential solutions. Plancraft addresses this gap by including a set of intentionally impossible tasks – tasks with no valid plan. This is similar to the task of identifying non-reachable atoms, as proposed in ACPBench (Kokel et al., 2025), or the unachievable tasks in WebArena (Zhou et al., 2023). By incorporating impossible tasks, Plancraft evaluates an agent’s capacity to assess task feasibility. This promotes more realistic and cost-effective decision making, encouraging agents to balance problem solving with estimates of feasibility.
@@ 3 Plancraft
To create Plancraft, we implement the logic and visual representation of the Minecraft crafting GUI. Like Prasad et al. (2023), we reconstruct the crafting process entirely in Python using game files and images from the Wiki: we use image manipulation to overlay items on top of the inventory background. This improves performance compared with using the Java game engine. Our environment has a one-to-one pixel mapping with the real Minecraft interface and supports multiple resolutions.
@@ 3.1 Assumptions
Plancraft makes a number of simplifying assumptions. First, there is a single agent in the environment, and all changes to the environment are a result of this agent’s actions. Second, the effects of executable actions are deterministic. Third, the crafting interface and inventory are fully observable, either through the text description or through the multi-modal image input. And finally, aligned with classical planning, actions are sequential: they are discrete and executed one at a time.
@@ 3.2 Action and Observation Space
The abstraction level chosen for the action and observation space impacts the tractability of the planning problem. Other Minecraft environments such as MineDojo (Fan et al., 2022) and Textcraft (Prasad et al., 2023) provide a high-level ‘craft’ command that doesn’t require the agent to manipulate the underlying ingredients. In contrast, Plancraft requires the correct placement of ingredients in the inventory and, as such, provides a lower-level symbolic action space. The two possible environment actions are smelt and move, and both actions require a slot_from, a slot_to and a quantity. This abstracts control dynamics and focuses on planning and spatial reasoning.
@@
In the Minecraft crafting GUI, there are $46$ possible item slots or positions. The crafting slot can be populated with an item only if the correct items have been deposited within the $3\times 3$ crafting grid—items can be moved in and out of this grid from the remaining 36 slots, in which items are stored. Items can only be withdrawn from the crafting slot, and doing this confirms the crafting operation and removes the items present in the crafting grid. Crafting recipes in Minecraft are either Shapeless or Shaped. Shapeless recipes only check whether the items required are present in the crafting grid, whereas Shaped recipes require the items to also be placed in the correct positions. When the items on the crafting grid match one of the $634$ recipes, then the resulting item is added to the crafting slot. We denote the crafting slot as [0], the $3\times 3$ crafting grid as slots [A1] to [C3] where the letter denotes each row and the number denotes the column, and the remaining 36 inventory slots as slots [I1] to [I36]. See Figure 2 for an example of a text description of the inventory that uses our annotation scheme.
Plancraft provides two types of observations: text-only and multi-modal. In text-only, agents receive a text description of the current inventory and crafting goal and the inventory is perfectly represented using our annotation scheme. Whereas in multi-modal, agents receive an image (dimensions $664\times 704$) that perfectly contains the crafting GUI, while the goal is provided as text.
@@ 3.3 Task Setup
To generate crafting tasks, we represent the dependencies among items as trees. The root of a tree is the target item while its leaves and intermediate nodes are the materials required to craft that item. For each crafting task, we sample the required recipes for the target item and then recursively explore the necessary materials to craft it. This builds a step-by-step plan from raw materials to the final product, simulating the crafting processes with varying levels of complexity. In addition to the necessary crafting materials, we also sample a number ($4$, $8$, $16$) of distractor items to make identifying the correct crafting path more challenging. Finally, we validate that all generated tasks can be solved in the text-only and multi-modal environments using a planner (see Section 3.4). We exclude invalid examples and ones where the handcrafted planner’s trajectory exceeds 30 steps.